OpenStudy (anonymous):

what are the dimensions of the rectangle below if the area is 100ft^2??

OpenStudy (anonymous):

@hartnn help!!!!!!!!

OpenStudy (tkhunny):

x * (x-15) = 100 where x > 0

OpenStudy (anonymous):

So you mean that x^2-15x would be the answer right?

OpenStudy (anonymous):

For the simpliefied expression that represents the perimeter of the rectangle right?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (anonymous):

I need helped bit time!!!! i have a paper to do & this is the only problem that i need!

OpenStudy (anonymous):

@electrokid ?

OpenStudy (mertsj):

Get it equal to 0 and factor it.

OpenStudy (mertsj):

x^2-15x-100=0

OpenStudy (anonymous):

So would it be x^2-15x=0 right then it would be x^2-15x-0 right?

OpenStudy (anonymous):

wait nevermind so then what 2 number could give 100 & 15?0_o i got 25 & 4 but that wont give me 15

OpenStudy (mertsj):

\[x^2-15x-100=(x-20)(x+5)\]

OpenStudy (anonymous):

so the expression that represents the perimeter of the rectangle is that right? & then for the leght would be 20 for x & for the width it would be 5 right?

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (anonymous):

@electrokid ?

OpenStudy (anonymous):

Is it good?

OpenStudy (mertsj):

Did you set each of those factors to 0?

OpenStudy (anonymous):

Nope. wait but its +20 & -5 right?

OpenStudy (mertsj):

And would -5 be appropriate for the side of a rectangle?

OpenStudy (anonymous):

i dont think so because its negative & i suppose to be positive & the area is positive so no?

OpenStudy (mertsj):

So x = 20 and the other side of the rectangle is x-15.

OpenStudy (anonymous):

so the width is 20 & the lenght is 5 right?

OpenStudy (mertsj):

I believe we would say the length is 20 and the width is 5. That is customary.

OpenStudy (anonymous):

wait so whats the simplified expression that represents the perimeter of the rectangle? was it x^2-15x-100?

OpenStudy (anonymous):

@Mertsj?

OpenStudy (mertsj):

How do you find perimeter?

OpenStudy (anonymous):

by adding but the question wants it in a simplifies form though so would it be 2x+2(x-15)?

OpenStudy (mertsj):

That is the expression but it is not simplified. Use the distributive property and add the like terms

OpenStudy (anonymous):

so would it be 2x+ 2x(1-15)? OR! 2X(1-15)?

OpenStudy (mertsj):

What is 2(x-15)

OpenStudy (anonymous):

The lenght

OpenStudy (mertsj):

Have you ever heard of the distributive property?

OpenStudy (anonymous):

i originally had x^2-15x but i dont know if thats right though?

OpenStudy (mertsj):